Find the annual intrest rate
Principal=$5000
Balance=$15,399.30
Time=15 years
Compounding=daily

Answer by stanbon(75887)

Find the annual interest rate
Principal=$5000
Balance=$15,399.30
Time=15 years
Compounding=daily
A=P(1+r/n)^(nt)
15,399.30 = 5000(1+r/365)^(365(15))
3.07986 = (1+r/365)^(5475)
Take the log of both sides to get:
log3.07986 = 5475log(1+r/365)
8.92294019X10^-5= log(1+r/365)
Take the anti-log of both sides to get:
1.00020548=1+r/365
r/365=0.00020548
r=0.07499998
Annual rate is 7.5%
Cheers,
Stan H.